OpenStudy (anonymous):

find the intervals on which f is increasing and decreasing \[\tan^{-1} x\]

OpenStudy (anonymous):

it is increasing on its domain

OpenStudy (anonymous):

Differentiate... i.e., Find the derivative of f(x)... It would never be lesser than 0...So always increasing...

OpenStudy (anonymous):

domain of \(\arctan(x)\) is \((-\frac{\pi}{2},\frac{\pi}{2})\) as \(x\) goes from \(-\frac{\pi}{2}\) to \(\frac{\pi}{2}\) \(\tan^{-1}(x)\) goes from \(-\infty\) to \(\infty\)